OpenStudy (anonymous):

any one know how to graph a line with slope -1 passing through the point (-2,-2)?

OpenStudy (anonymous):

get the equation for the line by using point slope formula, then draw a graph for that equation

OpenStudy (anonymous):

ok so I already have the slope which is y=1x + but I dont kow what to do with the (-2,-2)?

OpenStudy (anonymous):

-1x*

OpenStudy (anonymous):

y - y1 = m(x - x1) y - (-2) = -1[x-(-2)] y + 2 = -1[x + 2] y + 2 = -x - 2 y = -x - 2 -2 y = -x -4

OpenStudy (anonymous):

Now you can put x = -4 and calculate y for it then put x = 0 and calculate y for it finally put x = 3 and calculate y for it this will give you three pairs of x and y coordinates plot them and join them and u will hv your graph

OpenStudy (anonymous):

Do u understand ???

OpenStudy (anonymous):

No I need to see it in an equation like y=mx+b the lines must pass through -2,-2 idk im not getting it but thanks for your help I dont want to take up your time

OpenStudy (rosedewittbukater):

The slope is how many units the next point goes up or down divided by how many units the next point is to the side. Plug in the coordinates of the point to find the y intercept (b). y = mx + b y = -1x + b -2 = -1(-2) +b -2 = 2 + b -4 = b So the y intercept is -4, meaning that the line crosses over the horizontal axis at -4. The equation for the line is y=x - 4. To graph it, make a small x|y chart(or do it in your head) and plug in numbers like -2, -1, 0, 1, and 2 to get some points to graph a line. y = x-4 means that all of the y coordinates are going to be 4 less than their corresponding x's. x | y ----- -2| -6 -1|-5 0 |-4 1 |-3 2 |-2 So graph those coordinates of points. (-2, -6), (-1, -5), (0, -4), (1, -3), (2, -2) and then use a ruler or straight edge to draw a line through the points. Make sure to put little arrows at both ends of the line to imply that the line goes on forever. Hope this helped you understand how it works so you can do other problems like this. :)

OpenStudy (anonymous):

Oh my goodness that you soo much rosedewittbukater really really helpful Im going to write this down in my notes thank you so much

OpenStudy (rosedewittbukater):

Glad I could help.

OpenStudy (anonymous):

@rosedewittbukater @trinawhite31 rosedesitt.....your equation y = x-4 is wrong as it will not pass through the point (-2, -2) !!!! my method and solution is correct

OpenStudy (anonymous):

aww man is it wrong? you didnt give me any points like him thats why i didnt understand

OpenStudy (anonymous):

I explained it to you... my equation is y = -x -4 put x = -4, then y = -(-4) - 4 => y = 4 - 4 = 0 ---> gives point (-4, 0) now put x = 0, then y = -(0) - 4 => y = -4 ------> gives point (0, -4) finally put x = 3, then y = -(3) - 4 => y = -3 - 4 = -7 --> gives point (3, -7) now plot these point and you hv your graph.. Is it clear now???

OpenStudy (anonymous):

yes much more clear thank you

OpenStudy (anonymous):

you can chek that if you put x = -2 in my equation you get y = -(-2) - 4 = +2 - 4 = -2 and so you get the point (-2, -2) which has been given to you

OpenStudy (anonymous):

Yes I checked youre correct when I graphed it the line passed through (-2,2) thanks soo much Harkirat.

OpenStudy (anonymous):

u r welcome...☺

OpenStudy (rosedewittbukater):

Oh yes, Harkirat is correct. I accidentally changed the slope to 1 instead of -1 halfway through when I was typing. However the steps are the same, just change the equation to y = -x -4 and find the points the same way. Plug in some numbers to find a few points to make your graph. x | y ----- -2| -2 -1| -3 0 | -4 1 | -5 2 | -6 Then you plot the points (-2, -2), (-1, -3), (0, -4), (1, -5), (2, -6). Sorry for the misunderstanding, my mistake. The steps I gave you the first time remain the same, but the slope will change the equation of the graph.

OpenStudy (rosedewittbukater):

I hope I didn't confuse you too much. If you still decide to write the example in your notes, be sure to make the changes I did to have the equation y = -x -4. Basically after you find your equation that has the given slope, you insert some numbers into the equation to find the points like I did. For example, using -2 as an x value, substitute it for x. y = -1(-2) -4 y = 2-4 y = -2 That gives you your y value, which makes the point (-2, -2).

OpenStudy (anonymous):

Its ok every one makes mistakes and yes ive moved on to another question now.